(Question 17 of 26) A potential buyer wants to know about the biggest average increase in share price year on year. What was£70 £290 £310 £550 £860
Answer #1

\large \pounds 70

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

A B C D Change in Price Year 2014 2015 2016 2017 £ £ £ E Average share price 5.27 4.91 5.64 6.12 - £ € 0.36 0.73 0.48 biggest

Cell reference -

B Year 2014 Average share price Change in Price 5.27 =C4-C3 5.64 =C5-C4 6.12 =C6-C5 2015 2016 2017 4.91 biggest increase in a

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.